OpenStudy (anonymous):

if f(x)=3x^2-2x+1 and f(g(x))=12x^2-16x+6 now g(x)=?

OpenStudy (anonymous):

f(g)(x)=2(6x^2-8x+3)

OpenStudy (anonymous):

12x^2+6=f(g)(x)+16x

OpenStudy (anonymous):

no no f(x)=false!!!

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

f(g)(x)=false!!

OpenStudy (anonymous):

may you solve it?

OpenStudy (hoblos):

12x^2-16x+6 = 12x^2 - 12x +3 -4x +2 +1 = 3(4x^2 -4x+1) - 2(2x-1) +1 = 3(2x-1)^2 - 2(2x-1) +1 so g(x)= 2x-1

OpenStudy (anonymous):

oh yeah thanks

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):

Nicely done hoblos! 100 medals for you :D

OpenStudy (hoblos):

thanks ffm :D

OpenStudy (anonymous):

I don't know why I was thinking so much complicated :P KISS works here too :D

OpenStudy (anonymous):

KISS: keep it simple silly :P

OpenStudy (hoblos):

lol

OpenStudy (anonymous):

foolformath! Give me a kiss?